# Scalable Estimation of Multinomial Response Models

### v4

@@ Acknowledgment
$(\infty,5)$ $50$ 0.026 ( 0.15 ; 0.90 ) 0.211 ( 0.45 ; 0.94 ) 0.17 ( 0.45 ; 0.98 ) 0.107 ( 0.37 ; 0.98 ) 0.425 ( 0.03 ) 1.65
$200$ 0.004 ( 0.07 ; 0.98 ) 0.072 ( 0.23 ; 0.86 ) 0.07 ( 0.23 ; 0.90 ) 0.032 ( 0.18 ; 0.98 ) 0.266 ( 0.02 ) 5.73
$(\infty,15)$ $50$ 0.006 ( 0.08 ; 0.96 ) 0.073 ( 0.24 ; 0.92 ) 0.058 ( 0.24 ; 0.94 ) 0.041 ( 0.21 ; 0.96 ) 0.373 ( 0.02 ) 2.26
$200$ 0.002 ( 0.04 ; 0.94 ) 0.012 ( 0.12 ; 0.96 ) 0.011 ( 0.12 ; 0.96 ) 0.008 ( 0.1 ; 1.00 ) 0.201 ( 0.01 ) 7.64
$(1,5)$ $50$ 0.03 ( 0.15 ; 0.92 ) 0.507 ( 0.46 ; 0.78 ) 0.463 ( 0.46 ; 0.76 ) 0.138 ( 0.38 ; 0.98 ) 0.624 ( 0.03 ) 1.41
$200$ 0.005 ( 0.07 ; 0.96 ) 0.493 ( 0.24 ; 0.38 ) 0.488 ( 0.24 ; 0.38 ) 0.038 ( 0.19 ; 0.90 ) 0.548 ( 0.01 ) 4.68
$(1,15)$ $50$ 0.006 ( 0.08 ; 0.96 ) 0.097 ( 0.25 ; 0.88 ) 0.074 ( 0.24 ; 0.94 ) 0.042 ( 0.21 ; 0.94 ) 0.422 ( 0.02 ) 2
$200$ 0.002 ( 0.04 ; 0.92 ) 0.019 ( 0.12 ; 0.90 ) 0.019 ( 0.12 ; 0.94 ) 0.008 ( 0.1 ; 1.00 ) 0.277 ( 0.01 ) 6.38
  • For $\beta$ and $\delta$, we show the average of the mean squared errors (MSEs), using the posterior means as point estimators. In parentheses, the average posterior standard deviations (SDs) and empirical coverages of 95% credible intervals are shown. For $\pi$, we show the average of $L_{1}$ norm between the posterior mean and $\bm{\pi}^{*}$ with the posterior SDs in the parentheses. Time is the average seconds taken for sampling 1,000 MCMC draws in Matlab on a desktop with a 4.9GHz processor and 64GB RAM.
@@
The rest of the simulation design is the same as in Section 5.1 of the paper. Table SE.3 shows the simulation result. Crawford et al. (2021) show that incorrectly adding items to the consideration sets lead to biased estimates, which is reflected in the increased MSEs, especially when $n$ is large, compared to Table 2 of the paper where the time invariance assumption holds. In addition, there are distortions of the coverage.
We still see that the MSE/SD decrease in $n$ under the relatively large $T=15$. This is because the sample with $T=15$ has more periods with stable consideration sets than $T=5$. Thus, although the violation of the time-invariant consideration sets leads to issues known in the literature, and we are not an exception, our approach delivers reasonable result when there are enough periods during which the invariant assumption holds.
@@
$(K,T)$ $n$ $\beta$ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ Time
$(\infty,5)$ $50$ 0.032 ( 0.13 ; 0.8 ) 0.216 ( 0.33 ; 0.84 ) 0.151 ( 0.34 ; 0.92 ) 0.182 ( 0.34 ; 0.94 ) 1.59
$200$ 0.023 ( 0.06 ; 0.48 ) 0.196 ( 0.16 ; 0.3 ) 0.153 ( 0.17 ; 0.48 ) 0.046 ( 0.17 ; 0.9 ) 5.57
$(\infty,15)$ $50$ 0.01 ( 0.07 ; 0.82 ) 0.032 ( 0.15 ; 0.92 ) 0.039 ( 0.15 ; 0.88 ) 0.026 ( 0.18 ; 0.9 ) 2.15
$200$ 0.004 ( 0.04 ; 0.76 ) 0.016 ( 0.08 ; 0.7 ) 0.011 ( 0.08 ; 0.88 ) 0.009 ( 0.09 ; 0.96 ) 7.16
  • For $\beta$ and $\delta$, we show the average of the mean squared errors (MSEs), using the posterior means as point estimators. In parentheses, the average posterior standard deviations (SDs) and empirical coverages of 95% credible intervals are shown. For $\pi$, we show the average of $L_{1}$ norm between the posterior mean and $\bm{\pi}^{*}$ with the posterior SDs in the parentheses. Time is the average seconds taken for sampling 1,000 MCMC draws in Matlab on a desktop with a 4.9GHz processor and 64GB RAM.

### v5

@@ Acknowledgments
RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov L1-error (MCE) SD (ESD)       Cov
$(\infty,5)$ $50$ 0.15 ( 0.009 ) 0.15 ( 0.143 ) 0.96 0.473 ( 0.022 ) 0.46 ( 0.473 ) 0.94 0.437 ( 0.022 ) 0.46 ( 0.435 ) 0.96 0.341 ( 0.018 ) 0.36 ( 0.342 ) 0.97 0.434 ( 0.009 ) 0.03 ( 0.027 ) 0.97 1.84
$100$ 0.107 ( 0.006 ) 0.1 ( 0.103 ) 0.95 0.327 ( 0.017 ) 0.33 ( 0.315 ) 0.96 0.347 ( 0.02 ) 0.32 ( 0.334 ) 0.92 0.27 ( 0.013 ) 0.26 ( 0.269 ) 0.93 0.347 ( 0.005 ) 0.02 ( 0.021 ) 0.95 3.4
$(\infty,15)$ $50$ 0.082 ( 0.004 ) 0.08 ( 0.081 ) 0.96 0.252 ( 0.013 ) 0.24 ( 0.25 ) 0.93 0.243 ( 0.013 ) 0.24 ( 0.242 ) 0.96 0.204 ( 0.012 ) 0.2 ( 0.204 ) 0.95 0.368 ( 0.005 ) 0.03 ( 0.024 ) 0.96 2.58
$100$ 0.062 ( 0.003 ) 0.06 ( 0.059 ) 0.94 0.177 ( 0.01 ) 0.17 ( 0.175 ) 0.96 0.166 ( 0.008 ) 0.17 ( 0.164 ) 0.96 0.144 ( 0.008 ) 0.14 ( 0.143 ) 0.96 0.301 ( 0.003 ) 0.02 ( 0.018 ) 0.93 5.09
$(1,5)$ $50$ 0.155 ( 0.009 ) 0.15 ( 0.152 ) 0.95 0.673 ( 0.028 ) 0.48 ( 0.556 ) 0.82 0.675 ( 0.032 ) 0.48 ( 0.544 ) 0.84 0.364 ( 0.019 ) 0.37 ( 0.363 ) 0.97 0.77 ( 0.005 ) 0.03 ( 0.024 ) 0.78 1.56
$100$ 0.109 ( 0.006 ) 0.11 ( 0.109 ) 0.95 0.698 ( 0.029 ) 0.33 ( 0.413 ) 0.6 0.733 ( 0.031 ) 0.34 ( 0.456 ) 0.6 0.283 ( 0.014 ) 0.27 ( 0.28 ) 0.94 0.772 ( 0.005 ) 0.02 ( 0.019 ) 0.49 2.82
$(1,15)$ $50$ 0.085 ( 0.004 ) 0.08 ( 0.081 ) 0.96 0.292 ( 0.014 ) 0.25 ( 0.268 ) 0.9 0.281 ( 0.014 ) 0.25 ( 0.259 ) 0.92 0.208 ( 0.012 ) 0.2 ( 0.209 ) 0.96 0.711 ( 0.002 ) 0.02 ( 0.019 ) 0.89 2.27
$100$ 0.065 ( 0.004 ) 0.06 ( 0.06 ) 0.92 0.225 ( 0.012 ) 0.17 ( 0.192 ) 0.88 0.214 ( 0.01 ) 0.17 ( 0.182 ) 0.87 0.148 ( 0.008 ) 0.14 ( 0.145 ) 0.97 0.705 ( 0.001 ) 0.01 ( 0.014 ) 0.85 4.29
  • For $\beta$ and $\delta$, for each case, we show the estimated root mean squared error (RMSE), using the posterior means as point estimator. In parenthesis, the jackknife estimate of Monte Carlo Error (MCE) for the RMSE is presented. Next, the average of the posterior standard deviations (SD) is shown with the empirical standard deviation (ESD) of the posterior mean in the parenthesis. Third, the empirical coverage (Cov) of 95% credible interval is given.
  • For $\pi$, we show the average of $L_{1}$ norm between the posterior mean and $\bm{\pi}^{*}$ (L1-error). In the parenthesis, we show its jackknife estimate of MCE. The SDs, ESDs, and Covs are averaged over the 15 elements in $\pi$.
  • Time is the average seconds taken for sampling 1,000 MCMC draws in Matlab on a desktop with a 4.9GHz processor and 64GB RAM. The study is based on $R=200$ replications. 2,000 MCMC draws are obtained for each replication. The average of the inefficiency factors is around 6.8 with standard deviation 1.2.
@@
The rest of the simulation design is the same as in Section 5.1 of the paper. Table SE.3 shows the simulation result. Crawford et al. (2021) show that incorrectly adding items to the consideration sets lead to biased estimates, which is reflected in the increased RMSEs, especially when $n$ is large, compared to Table 2 of the paper where the time invariance assumption holds. In addition, there are distortions of the coverage.
We still see that the RMSEs/L1-errors/SDs clearly decrease in $n$ under the relatively large $T=15$. This is because the sample with $T=15$ has more periods with stable consideration sets than $T=5$. Thus, although the violation of the time-invariant consideration sets leads to issues known in the literature, and we are not an exception, our approach delivers reasonable result when there are enough periods during which the invariant assumption holds.
@@
$(K,T)$ $n$ $\beta$ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ $\pi$ Time
RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov L1-error (MCE) SD (ESD)       Cov
$(\infty,5)$ $50$ 0.184 ( 0.008 ) 0.13 ( 0.143 ) 0.81 0.483 ( 0.02 ) 0.33 ( 0.366 ) 0.78 0.387 ( 0.019 ) 0.34 ( 0.324 ) 0.91 0.407 ( 0.025 ) 0.34 ( 0.373 ) 0.92 0.724 ( 0.012 ) 0.03 ( 0.029 ) 0.9 1.78
$100$ 0.163 ( 0.006 ) 0.09 ( 0.095 ) 0.67 0.409 ( 0.015 ) 0.23 ( 0.227 ) 0.69 0.364 ( 0.014 ) 0.24 ( 0.218 ) 0.8 0.29 ( 0.014 ) 0.24 ( 0.23 ) 0.91 0.688 ( 0.008 ) 0.02 ( 0.022 ) 0.94 3.33
$(\infty,15)$ $50$ 0.091 ( 0.004 ) 0.07 ( 0.077 ) 0.91 0.178 ( 0.008 ) 0.15 ( 0.16 ) 0.89 0.167 ( 0.008 ) 0.16 ( 0.154 ) 0.94 0.173 ( 0.01 ) 0.18 ( 0.164 ) 0.93 0.707 ( 0.009 ) 0.02 ( 0.024 ) 0.87 2.51
$100$ 0.073 ( 0.003 ) 0.05 ( 0.055 ) 0.84 0.143 ( 0.006 ) 0.11 ( 0.108 ) 0.86 0.134 ( 0.005 ) 0.11 ( 0.105 ) 0.88 0.116 ( 0.005 ) 0.12 ( 0.112 ) 0.97 0.656 ( 0.006 ) 0.02 ( 0.018 ) 0.93 4.79
  • For $\beta$ and $\delta$, for each case, we show the estimated root mean squared error (RMSE), using the posterior means as point estimator. In parenthesis, the jackknife estimate of Monte Carlo Error (MCE) for the RMSE is presented. Next, the average of the posterior standard deviations (SD) is shown with the empirical standard deviation (ESD) of the posterior mean in the parenthesis. Third, the empirical coverage (Cov) of 95% credible interval is given.
  • For $\pi$, we show the average of $L_{1}$ norm between the posterior mean and $\bm{\pi}^{*}$ (L1-error). In the parenthesis, we show its jackknife estimate of MCE. The SDs, ESDs, and Covs are averaged over the 15 elements in $\pi$.
  • Time is the average seconds taken for sampling 1,000 MCMC draws in Matlab on a desktop with a 4.9GHz processor and 64GB RAM. The study is based on $R=200$ replications. 2,000 MCMC draws are obtained for each replication. The average of the inefficiency factors is around 6.4 with standard deviation 1.1.
SE.5 Additional simulation results for $J=100$
(a) $T=1$
(b) $T=5$
(c) $T=200$
Figure SE.4: Similarity matrices. The results are based on one replication of simulated data with $J=100$, $T=\{1,5,200\}$, and $n=100$. In the true clustering, the first subpopulation contains the units 11 to $50$ and the second contains the remaining units.
In Section 5.2, we present the simulation results for the high dimensional case with $J=100$. Although it is not possible to show the results for $\bm{\pi}$ due to the $2^{100}-1$ support points, we illustrate its estimation result in Figure SE.4 that shows the $n\times n$ “similarity matrices” based on one replicate of the simulation data. These give the posterior probability that a given subject in a particular row $k$ is in the same cluster as another subject at a specific column $i$ which is computed as the posterior probability of the event $\{S_{k}=S_{i}\}$. This probability ranges from zero (light blue) to one (dark blue). By $T=5$, the similarity matrix roughly aligns with the true clustering structure, leading to accurate estimates of $\bm{\pi}$. For $T=200$, the clustering structure is recovered with high accuracy.